OpenStudy (anonymous):

Suppose "p" varies directly as "d", and p=2 when d=7. What is the value of "d" when p=10? A. d=20/7 B. d=35 C. d=7/5 D. d=15 ~Thanks ;)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

"p" varies directly as "d" means that p = k*d

jimthompson5910 (jim_thompson5910):

p = 2 when d = 7 p = k*d 2 = k*7 k = ???

OpenStudy (anonymous):

Still confused... I don't get what the deal is with k

jimthompson5910 (jim_thompson5910):

k is just a constant

jimthompson5910 (jim_thompson5910):

it's some fixed number

jimthompson5910 (jim_thompson5910):

if 2 = k*7, then k must be ???

OpenStudy (anonymous):

Ohhh... 2/7

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

p = k*d p =(2/7)*d 10 = (2/7)*d ... plug in p = 10 now solve for d

OpenStudy (anonymous):

B. d=35 Thanks! Was not as difficult as I was making it.

jimthompson5910 (jim_thompson5910):

yeah it's not too bad once you get the hang of it